OpenStudy (anonymous):

Determine whether the polynomial is a perfect square and if it is, factor it. n^2 – 9n + 9 A. is a perfect square: (n – 3)^2 B. is a perfect square: (n – 9)^2 C. is not a perfect square D. is a perfect square: (n + 3)^2

OpenStudy (anonymous):

go to www.wolframalpha.com and maybe ull find ur answer it shows a bunch of data for that

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

When you are factoring something of the form: \[x^2 +ax + b^2\] (like what you have here), it'll be a perfect square if a = 2b Another way to think about this is, given the polynomial you have, can you find a number that, when squared, will equal the number with no x on it (in this case, 9), and when added to itself (or multiplied by 2), will equal the middle number with just x on it (in this case, -9) If you can, it's a perfect square. For instance: \[x^2 +8x + 16\] I need a number that when squared will equal 16, and when multiplied by 2 will equal 8. That number is 4, so: \[x^2+8x+16 = (x+4)^2\] Can you find such a number for your polynomial?

OpenStudy (anonymous):

\[\left( a+b \right)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

\[\left( a-b \right)^2=a^2-2ab+b^2\]

OpenStudy (campbell_st):

well start by looking at 9 \[9 = (\pm 3)^2\] and \[x^2 = (\pm x)^2\] so does the middle term have a value of \[2 \times (\pm 3) \times (\pm x)\] if it does, you have a perfect square... if it doesn't its not a perfect square

OpenStudy (anonymous):

Im so confused..

OpenStudy (anonymous):

It was C -.-

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I just randomly guessed and got it right ^.^

OpenStudy (anonymous):

nice!